the average wave diary wavelength of the yellow light emitted from a sodium lamp is 5893 angstroms Express it in NM

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
As you should know
 
1 metre = 10^10 A°
&
1 metre = 10^9 Nano metres
Hence
 
1 n m = 10 A ° 
Or
1 Angstroms (A°) = 1/10  nm
Hence
5893 A° = 589.3 nm
 
Regards
Arun (askIITians forum expert)